# Keyed Chaotic Masking: A Functional Privacy Framework for Neural Inference

### v2

@@ 3.3 Chaos Diagnostics and Characterization
To ensure that the graph dynamical systems constructed using cryptographic keys exhibit chaotic behavior, we perform a series of diagnostics drawn from nonlinear systems theory. Chaos is typically defined by the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small differences in initial conditions. This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories, providing both privacy and resistance to inversion.
3.3.1 Lyapunov Exponent Estimation.
@@
3.4 Key-Conditioned System Assembly
The final stage of the construction process assembles the graph dynamical system by combining the sampled topology, selected dynamical rule, and initial state—all deterministically derived from a shared cryptographic key. This key serves as the sole source of entropy for the transformation and ensures that the system is uniquely bound to the key holder.
To maintain predictable chaotic behavior, the method restricts sampling to a curated set of parameter ranges known to yield high sensitivity and non-periodicity. Each candidate configuration includes:
  • A graph family (e.g., Erdős–Rényi, Watts–Strogatz)
  • Associated hyperparameters (e.g., edge probability $p$, rewiring rate $\beta$, node degree $k$)
  • A dynamical rule (e.g., logistic map, Izhikevich system)
  • Rule-specific parameters (e.g., $r$ for logistic map, coupling strength $\gamma$, integration constants)
These configurations are indexed, and a pseudorandom generator seeded from the key selects from this indexed set. This ensures that the resulting system is both deterministic and cryptographically unique, while also avoiding regimes where dynamics are trivial or periodic.
In addition to the structural parameters, the initial state of the system is randomly sampled using the same seed derived from the key. This means both the system’s structure and starting conditions are key-dependent, ensuring that even if two users select the same dynamical configuration, their trajectories will diverge due to differing initial conditions.
The result is a fully assembled, chaotic transformation system that is:
  • Deterministic: reproducible with the key
  • Key-obscured: opaque to observers without the key
  • Composable: the system can be integrated into existing pipelines without modifying model architecture or training objectives
This construction enables lightweight, cryptographically bound transformations.
@@ 4 Applications
The applications described in this section are conceptual demonstrations based on the properties of the transformation system. While they are technically implementable and initial implementation confirms Lyapunov divergence across key configurations; empirical evaluations of inference compatibility, robustness to key error, and resistance to inversion attacks are ongoing. In all applications, we assume a shared key is either held by the user or securely negotiated with the server via key exchange (e.g., ECDH). The transformation can be applied on either side of the interaction, depending on the privacy model and deployment scenario.
@@ 4.1 Key-Authenticated Neural Network Inference
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is masked client-side via:
@@
Optionally, the model can also apply a forward scrambling to its outputs using a second chaotic trajectory:
@@
This creates a fully authenticated inference pipeline, where both input and output representations are cryptographically tied to the user’s key. This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect data in transit.
4.2 Watermarking for Model & User Attribution
In multi-user inference systems, it is often desirable to verify which user or device generated a particular output—especially in contexts involving auditing, content moderation, or abuse prevention. While some systems incorporate output watermarking to identify model-generated content, these are typically model-wide and do not support per-user attribution.
The proposed framework enables per-user output watermarking by applying a deterministic, key-conditioned transformation to the model output. Let $\mathbf{y}\in\mathbb{R}^{d}$ represent the output of a model (e.g., a logit vector, embedding, or generated content representation). A user-specific watermark is applied as:
$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
where $\mathbf{S}^{\prime}_{k}\in\mathbb{R}^{d}$ is a deterministic mask generated from the user’s key using the same chaotic transformation framework. This watermark is embedded server-side before the output is returned to the user or downstream system.
Crucially, this transformation acts as a soft fingerprint: the output remains usable for its original task, but contains a key-specific signal that can later be used to attribute the output to a specific user. If a model output appears in the wild, the server or auditor can regenerate $\mathbf{S}^{\prime}_{k}$ for known keys and match it against the suspicious output to infer its origin.
This watermarking mechanism enables per-user attribution without modifying the model, embedding user IDs, or relying on metadata. It offers a lightweight path to traceable outputs in shared inference environments, content moderation systems, or multi-tenant deployment pipelines.
4.3 Data Privacy
Many inference systems operate on embeddings derived from sensitive user data, such as personal text, biometric signals, or proprietary representations. In these settings, raw embeddings are often transmitted to centralized models for downstream inference—posing a significant privacy risk if intercepted or logged. While full encryption can provide protection, it typically requires secure execution environments or homomorphic computation, which introduces significant overhead.
The proposed transformation scheme offers a lightweight alternative by obfuscating embeddings directly in their native vector space. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of token or patch embeddings. The user applies a chaotic, key-conditioned transformation as follows:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$. The masked embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically transformed and difficult to interpret without knowledge of the system.
Importantly, because the transformation is deterministic and reversible with the key, the model hosted on the server can decode the input by regenerating the same mask internally:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead. Furthermore, reverse engineering the obscured tensors is impractical without access to the full chaotic transformation, as the dynamics induce exponential divergence with respect to both key and input, making inversion intractable without the key.
4.4 Secure Data Contribution to Shared Models
In collaborative machine learning systems—such as federated learning, crowd-sourced training, or multi-institutional data sharing—users may wish to contribute data to a central model while retaining control over their raw inputs. However, direct data sharing raises concerns about privacy, misattribution, and unauthorized reuse, particularly in domains involving personal or proprietary information.
The proposed transformation framework allows users to submit data masked by a key-conditioned chaotic transformation. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$. The transformed data is computed as:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
and contributed to a shared model or data repository. The transformation preserves structure and dimensionality, but obfuscates semantic content, thereby protecting raw data from inspection or unauthorized reuse.
In some scenarios, the server may invert the transformation:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
to recover the original data for training. Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer. However, training directly on transformed data may degrade performance depending on the task, and its effectiveness is likely to be application-specific. Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
@@ 5 Discussion
This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to provide input obfuscation, output watermarking, and authenticated inference in modern machine learning pipelines. The system is lightweight, deterministic, and modular—requiring no changes to model architecture.
While the system is designed to offer functional security properties such as key-conditioned masking and trajectory-based unpredictability, these remain to be empirically evaluated in practical deployments. In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models. As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
The choice of dynamical system (e.g., graph topology, update rule, initialization) plays a significant role in determining the transformation’s properties. While this work restricts itself to a curated set of configurations with empirically chaotic behavior, future work could explore learned dynamics, adversarial training of transformation functions, or differentiable chaos-conditioned layers for integration into model backbones.
More broadly, this work suggests that chaotic systems—often studied in the context of memory, control, or biological modeling—can be repurposed as lightweight cryptographic primitives for AI deployment. As the interface between private data and hosted models becomes increasingly important, new methods that offer user-level control, attribution, and secure signaling will be essential. This framework provides one such direction and opens several avenues for further development.
@@ 6 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model. The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential. As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation. Please note that this preprint version is released to share the concept and invite community feedback. Validation experiments and security evaluations are ongoing and will appear in subsequent revisions.
@@ 7 Acknowledgements
The author gratefully acknowledges the UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas. This work was completed independently during a period of academic leave. The author also acknowledges the use of large language models (LLMs) to support drafting, editing, and idea development during the research and writing process.

### v3

@@ 3.2.3 Chaotic Dynamics Diagnostics
To ensure that the graph dynamical systems constructed using cryptographic keys exhibit chaotic behavior, we perform a series of diagnostics drawn from nonlinear systems theory. Chaos is typically defined by the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small differences in initial conditions. This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories an important property for ensuring the security of this encrpytion method.
@@
3.3 Encryption Algorithm
This section consolidates the components described above into a complete encryption and decryption procedure. Given a cryptographic key, a deterministic chaotic graph dynamical system is instantiated by deriving subkeys and sampling graph structure and node dynamics. Forward dynamics is then simulated to produce a structured mask that is applied to the input tensor via element-wise addition in the case of encryption and subtraction in the case of decryption. The resulting process is lightweight, reproducible, and key-dependent—offering functional privacy, it can also be incorporated directly into existing neural architectures as a layer without requiring retraining of the model architecture. The full algorithm is outlined below:
Input: Input tensor $X$, secret key $k$, public nonce $n$
Output: Masked tensor $\tilde{X}$ or recovered tensor $X$
Derive subkeys $k_{\text{graph}},k_{\text{params}},k_{\text{init}},k_{\text{noise}}$ using SHA-256:;
foreach label $i\in\{\texttt{graph},\texttt{params},\texttt{init},\texttt{noise}\}$ do
       $k_{i}\leftarrow\text{SHA-256}(\texttt{label}_{i}\,\|\,k\,\|\,n)$;
Construct chaotic graph dynamical system:;
Sample graph topology $G$ from $k_{\text{graph}}$;
Configure parameters and initialize state from $k_{\text{params}},k_{\text{init}}$;
Seed a pseudorandom number generator with $k_{\text{noise}}$ to generate deterministic noise;
for each timestep $t=1$ to $T$ do
       Apply node updates using local dynamics and diffusive coupling;
       Inject deterministic noise into each node state at timestep $t$;
Aggregate final node states into mask matrix $S\in\mathbb{R}^{n\times d}$;
Encryption: $\tilde{X}\leftarrow X+S$;
Decryption: $X\leftarrow\tilde{X}-S$;
Algorithm 1 Keyed Chaotic Dynamics (Encryption and Decryption with Deterministic Noise)
@@ 4 Applications
This section outlines practical use cases enabled by the proposed transformation framework. While experimental validation is ongoing, the design properties of the system—key-conditioned masking, deterministic reversibility, and compatibility with real-valued tensor representations—make it suitable for deployment in a range of privacy-sensitive machine learning settings. In all cases, we assume a shared secret key $k$ is securely negotiated between parties (e.g., via ECDH) and that a public nonce $n$ is used to ensure per-input uniqueness.
@@ 4.1 Key-Authenticated Neural Network Inference
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic graph dynamical system from a cryptographic key $k$ and nonce $n$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is encrypted client-side via:
@@
Optionally, the model can also encrypt its outputs using a second chaotic trajectory:
@@
This creates a fully authenticated inference pipeline, where both input and output representations are cryptographically tied to the user’s key. This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect both data in transit and storage.
4.2 Privacy-Preserving Data Sharing and Inference
In both inference and collaborative learning settings, real-valued embeddings may contain sensitive information (e.g., personal text, biometrics, or proprietary features). The proposed transformation enables users to encrypt these representations locally using a key-conditioned chaotic system before transmitting them to a server or aggregator.
The encrypted input $\tilde{X}=X+S_{k}$ retains the structure and shape required for downstream tasks. Because the transformation is deterministic and reversible with the key and nonce, the server can reconstruct the original input for inference or training, this is especially relevant in TEE environments where the computational graph representing the encryption/decryption algorithm is not easily interpreted.
This enables secure, per-user control over data contributions—offering lightweight protection against inspection, misuse, or unauthorized reuse.
5 Security Considerations
5.1 Threat Model
This framework is designed as a lightweight privacy mechanism for inference-time data obfuscation and input authentication. It aims to:
  • Prevent unauthorized users from generating valid inputs to a model without access to a cryptographic key.
  • Obfuscate real-valued inputs during transit or storage without modifying model architecture.
  • Enable reversible masking for authorized parties in possession of the key.
We assume a standard threat model where:
  • The shared key $k$ is negotiated securely (e.g., via Elliptic Curve Diffie-Hellman).
  • The attacker has access to the masked tensor $\tilde{X}$, but not the key or the internal configuration of the mask-generating chaotic system.
  • The system is not used in repeated or low-entropy input settings without additional safeguards.
5.2 Security Properties
The proposed transformation offers the following functional guarantees:
  • Key- and nonce-conditioned obfuscation: The transformation from $X$ to $\tilde{X}$ is deterministic and reversible only with the correct cryptographic key $k$ and associated nonce $n$. This ensures unique masks per input and prevents mask reuse attacks, even when the same key is reused across queries.
  • Input authentication: Only users in possession of the correct key and nonce can produce valid masked inputs that result in successful model inference, effectively binding access to possession of the correct credentials.
  • Nonlinear unpredictability: The chaotic update rule introduces exponential sensitivity to both key and input perturbations. The injection of deterministic, key-derived noise at each timestep further increases trajectory divergence, degrading the feasibility of surrogate inversion or pattern learning.
  • Model compatibility: The transformation operates entirely in continuous space and is represented by a computation graph. It integrates seamlessly into existing deep learning pipelines and can be implemented efficiently on hardware accelerators such as GPUs using CUDA kernels or fused operations, enabling low-latency deployment even in real-time inference settings.
5.2.1 Implementation Considerations for Secure Deployment
The transformation can be efficiently implemented as a fused CUDA kernel to minimize memory access overhead and latency. However, this design choice introduces additional considerations: GPU memory and execution contexts are shared across processes, and on systems lacking strict memory isolation, adversaries with privileged access may attempt to inspect device memory or leverage side-channel leakage (e.g., timing or power-based signals) to infer key-conditioned behavior. While such attacks require elevated privileges and are not typically feasible in hosted model settings, developers deploying this method in untrusted or multi-tenant environments should consider GPU sandboxing, memory zeroing, and kernel isolation practices. Integrating the transformation within a trusted execution environment (TEE) or extending it with secure enclaving techniques may offer additional protection against kernel-level inspection.
5.3 Limitations and Attack Vectors
This transformation is best viewed as a functional privacy primitive for continuous tensor data—analogous in structure to symmetric encryption schemes such as AES, but adapted for real-valued, neural architectures. While it does not aim to replace formal cryptographic methods in settings requiring ciphertext indistinguishability, it offers a complementary security protocol for scenarios involving neural inference. Potential vulnerabilities include:
  • Model inversion attacks (unlikely): An adversary with full access to masked inputs and model outputs could theoretically attempt to learn a surrogate inverse mapping. However, the transformation is highly nonlinear, key-conditioned, and exhibits exponential sensitivity to perturbations via chaotic dynamics. Additionally, the injection of deterministic, per-timestep noise renders the resulting mask trajectories resistant to approximation, making such attacks computationally infeasible in practice.
  • Key compromise: The transformation is fully reversible by design. If the cryptographic key and associated nonce are leaked or mishandled, the original input can be trivially recovered by an adversary. Standard key management practices (e.g., ephemeral session keys, ECDH negotiation) should be used to mitigate this risk.
  • Lack of formal obfuscation guarantees: The transformation depending choice of dynamics applies a bijective, key-conditioned nonlinear mapping over continuous, real-valued tensors. While the resulting representations are semantically obfuscated and reversible only by authorized parties, the system does not rely on finite-field cryptographic primitives and does not satisfy formal definitions of indistinguishability. As such, it should be viewed as a structured, functional privacy mechanism rather than a substitute for conventional encryption in adversarial threat models.
5.4 Integration with Secure Execution
This system is intended as a complementary privacy layer. In high-risk deployments, it can be combined with:
  • Standard transport-layer encryption (e.g., TLS)
  • Trusted Execution Environments
  • Ephemeral key derivation or key rotation schemes
  • Obfuscation techniques such as dropout, quantization, or learned masking
5.5 Future Directions
Future work includes:
  • Formal analysis of the system’s resistance to statistical and adversarial attacks.
  • Investigation of dynamic or learned chaotic systems for adaptive masking.
  • Exploration of surrogate model defenses and integration into split-learning architectures.
This framework is best viewed as a flexible privacy primitive that enables data-level masking and authentication with minimal architectural assumptions. It is particularly useful in scenarios where full encryption is infeasible or unnecessary, but lightweight, key-bound transformations are desirable.
@@ 6 Discussion
This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to apply key-conditioned, reversible obfuscation to real-valued tensors within modern machine learning pipelines. The system is lightweight, deterministic, and modular—requiring no changes to model architecture and operating natively in continuous space. It is amenable to efficient GPU execution, including fused CUDA kernel implementations for real-time inference masking.
The transformation provides functional security guarantees such as cryptographic key- and nonce-conditioned masking, semantic obfuscation, and trajectory-level unpredictability. It supports both input authentication and reversible transformation, enabling inference-time privacy and secure data contribution without disrupting model compatibility. While the system is reversible by design for authorized users, its behavior in adversarial settings remains to be evaluated. In particular, the method does not rely on formal encryption primitives over finite fields and may be vulnerable to misuse if keys are leaked or mask reuse occurs without appropriate nonce management. As such, the transformation is best viewed as a complementary privacy layer—positioned between full encryption and raw data sharing—and can be combined with techniques such as differential privacy, trusted execution environments, or encrypted transport protocols in high-assurance contexts.
The design of the chaotic system (e.g., graph topology, local map, coupling strategy) plays a critical role in shaping both the security and representational properties of the transformation. This work restricts itself to a curated set of update rules with empirically verified chaotic behavior, but future directions include optimizing or learning system configurations from data, developing differentiable chaos layers, or adversarially training transformation functions to withstand surrogate model attacks.
More broadly, this work suggests that chaotic dynamical systems—traditionally studied in the context of memory, control, or biological computation—can be reimagined as lightweight, cryptographically inspired building blocks for AI systems. As machine learning continues to interface with sensitive and user-controlled data, techniques that support secure signaling, data attribution, and behavioral integrity will be increasingly essential. This framework provides one such path forward, and opens several directions for continued exploration.
@@ 7 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys and public nonces, the method enables encryption/decryption and authenticated inference based on keys without requiring retraining or alterations to original model architectures. The transformation operates directly on real-valued tensors in continuous space, making it compatible with modern neural pipelines and efficient to implement on hardware accelerators.
The proposed system serves as a functional privacy mechanism: key- and nonce-conditioned, reversible, and structurally opaque to unauthorized users. It offers a lightweight, modular alternative to traditional cryptographic methods for use cases involving secure inference, controlled data contribution, and user-authenticated API access.
As machine learning systems increasingly interact with sensitive and user-owned data, new methods that enable secure, interpretable, and user-aligned processing will become essential. This framework contributes one such approach, grounded in dynamical systems theory and designed for integration into existing model infrastructure. Validation experiments and empirical evaluations are ongoing. This preprint is shared to disseminate the concept, gather community feedback, and guide future research into learned chaotic dynamics, adversarial robustness, and secure deployment practices.
@@ 8 Acknowledgements
The author gratefully acknowledges their UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas. This work was completed independently during a period of academic leave. The author also acknowledges the use of large language models (LLMs) to support drafting, editing, and idea development during the research and writing process.